Circular No. 7551

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000fq D. M. Wittman, A. C. Becker, V. E. Margoniner, and J. A. Tyson, Bell

Laboratories, Lucent Technologies; I. P. Dell’Antonio, Brown University; D. Loomba, University of New Mexico; and R. A. Schommer, Cerro Tololo Interamerican Observatory (CTIO), report the discovery of a supernova (R ∼ 23.8) in a series of five R-band images taken on Dec. 21.4 UT with the 4-m Mayall telescope at Kitt Peak; the images were differenced with deeper (limiting R mag 26) images taken one year ago using the same instrumentation. SN 2000fq is located at α = 9^{h}23^{m}37^{s}.68, δ = +29^{o}52′26′′.0 (equinox 2000.0), which is > 5′′ from the nearest member of a group of galaxies (R ∼ 20). Images taken on Dec. 21.3 with the Sloan z filter also show the new object at z′ ∼ 22.6. SN 2000fq was reconfirmed by R. C. Smith and R. A. Schommer on Dec. 25.3 at R ∼ 24.0 using the Blanco 4-m telescope (+ Mosaic2 camera), and on Dec. 28.45 by H. Aussel and D. Sanders using the Canada-France-Hawaii telescope (+ UH12K mosaic) at R ∼ 24.2. These magnitudes are based on R = 20.90 for the star located 7′′.4 south of the supernova. Spectroscopic follow-up was obtained by D. Stern, A. Dey, S. Dawson, and H. Spinrad using the Keck-1 telescope on Dec. 24.5. The spectrum is consistent with that of a type-Ia supernova at a redshift of 0.7. Two of the galaxies in the group, to the east, are also at z = 0.7, with strong [O II] and a 400.0-nm break.

SUPERNOVA 1999by IN NGC 2841 R. Kirshner and P. Challis, Harvard-Smithsonian Center for Astro-

physics; P. Garnavich, University of Notre Dame; and the Supernova Intensive Study collaboration report that Hubble Space Telescope Wide Field/ Planetary Camera-2 (WFPC2) images of the peculiar type-Ia SN 1999by were obtained Sept. 28 with the F555W filter and again on Dec. 10 in both the F555W and F439W filters: “The derived magnitude 580 days after maximum is V = 24.6 ± 0.1, which when combined with earlier groundbased photometry indicates a decline rate in V of 0.0106 mag/day (250–600 days after maximum). This is significantly slower than the 0.0139 mag/day rate of SN 1992A (N. Suntzeff, private comm.) over the same interval after maximum light. The B-band decline rate of 0.0099 mag/day is also slow compared to that (0.0128 mag/day) for SN 1992A. The B − V color at day 580 is −0.4 ± 0.2 and is extremely blue for a type-Ia event, which typically show B − V = 0 at late times (Milne et al. 1999, Ap.J. Suppl. 124, 503).”

2000 December 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green